likeabossssssss (likeabossssssss):

A train travels 42 kilometers in 15 minutes. How far will the train travel in 3 hours at that speed? A. 168 km B. 336 km C. 504 km D. 672 km b?

likeabossssssss (likeabossssssss):

@mathmale

OpenStudy (mathmale):

I'd suggest you take the given info and write it as a "rate of change:"\[\frac{ 42~km }{ 15~\min }\]

OpenStudy (xguardians):

3 hours = 180 minutes 180 minutes / 15 minute intervals is 12 12 intervals * 42 = 504 C

OpenStudy (mathmale):

You could either use this rate as is, or you could reduce the numerical fraction. To find the distance traveled in 3 hours, convert 3 hours to minutes, and then multiply the rate by that number of minutes. Or, convert 15 min to hours and multiply the resulting rate by 3 hours. Please show your work.

likeabossssssss (likeabossssssss):

so 60/15=4 4*3=12 is that how u do it ?

likeabossssssss (likeabossssssss):

@mathmale

OpenStudy (xguardians):

@likeabossssssss look at my work, yes that is correct, then you multiply that by 42km

likeabossssssss (likeabossssssss):

it would be 540km

OpenStudy (mathmale):

In making the conversions we have to express the results in the proper units of measurement.\[\frac{ 42~km }{ 15~\min }*\frac{ 60~minutes }{ 1~hour }\]

likeabossssssss (likeabossssssss):

ok

OpenStudy (mathmale):

Then multiply this result by 3 hours. Ensure that the final units of measurement are Km.

likeabossssssss (likeabossssssss):

4*3= 12 then times it by 42 like @xGuardians said?

OpenStudy (mathmale):

Again, please show your units of measurement; without them, it's considerably harder to see what you're attempting. \[\frac{ 42~km }{ 15~\min }*\frac{ 60~minutes }{ 1~hour }*3~hours\]

likeabossssssss (likeabossssssss):

how do I show them

OpenStudy (mathmale):

I've modeled use of the units of measurement as best I can in the work shared with you (above).

likeabossssssss (likeabossssssss):

so 42km over 15mins times 60 mins over 1 hour times three?

OpenStudy (mathmale):

Yes. it'd be better to "do the math" with the formula work I've given you instead of changing the problem back to words. Start by cancelling units where possible; for example, cancel "minutes." Again, please show your work and stick with formulas / expressions.